Samsung currently manufactures its S Pen for the Galaxy Note phone series. This accessory has almost reached maturity, and the company is considering whether to replace it with a newer model S Pen stylus including a Bluetooth feature for their latest mobile Galaxy Note 9. It is projected that this newer item will sell for $30 each.

                               

The current model sells 75,000 units for $25 each with its manufacturing Process A requiring an investment of $120,000 for fixed costs, and results in a $19 per unit cost.

Technologies for this accessory have not changed drastically, with additional Bluetooth connectivity. Moreover, if it were to introduce the new S Pen stylus, Samsung is trying to decide whether to make or buy this accessory.

If the item ends up being outsourced, there is virtually no cost other than the $26 per unit that they would pay their supplier.

If the new item is manufacturer in hour, they have to switch their manufacturing process to Process B. This requires an investment of $150,000 for design and equipment, but results in a $24 per unit cost. Regardless of whether the item is subcontracted or produced internally, the company feels that the market has a 50 percent chance of being “receptive” to the new S Pen stylus and will sell 100,000 units. On the other hand, the market has a 30 percent chance of being “neutral” with sales of 75,000 units and 20 percent chance of being “unfavorable” with sales of 50,000 units.

Using decision trees and EMV, what is their best choice?

              

Samsung currently manufactures its S Pen for the Galaxy Note phone series. This accessory has almost reached maturity, and the company is considering whether to replace it with a newer model S Pen stylus including a Bluetooth feature for their latest mobile Galaxy Note 9. It is projected that this newer item will sell for $30 each.

                               

The current model sells 75,000 units for $25 each with its manufacturing Process A requiring an investment of $120,000 for fixed costs, and results in a $19 per unit cost.

Technologies for this accessory have not changed drastically, with additional Bluetooth connectivity. Moreover, if it were to introduce the new S Pen stylus, Samsung is trying to decide whether to make or buy this accessory.

If the item ends up being outsourced, there is virtually no cost other than the $26 per unit that they would pay their supplier.

If the new item is manufacturer in hour, they have to switch their manufacturing process to Process B. This requires an investment of $150,000 for design and equipment, but results in a $24 per unit cost. Regardless of whether the item is subcontracted or produced internally, the company feels that the market has a 50 percent chance of being “receptive” to the new S Pen stylus and will sell 100,000 units. On the other hand, the market has a 30 percent chance of being “neutral” with sales of 75,000 units and 20 percent chance of being “unfavorable” with sales of 50,000 units.

Using decision trees and EMV, what is their best choice?
Answer #1

Profit with current model = 75000x(25-19) -120000 =330000

If the company chooses to manufacture the new pen it has two options

(a) Outsource. In this case Profit will be = ( 30-26) xN  

where N = 0.5x100000+0.3x75000+0.2x50000 =82500

Profit with outsourcing = 82500x4 = 330000

(b) Make inhouse = (30-24)x82500 - 150000 =345000

Since the EMV is highest with making the new product inhouse, it is the right decision,